OpenStudy (anonymous):

sigma notation : upper limit is 6, lower limit is k=4.

OpenStudy (anonymous):

\[\sum_{k=4}^{6} (k-4)x= 450\]

OpenStudy (anonymous):

what the value of x?

OpenStudy (anonymous):

is it 150?

OpenStudy (anonymous):

X = 450 / sum = 450/ 2 = 225

OpenStudy (anonymous):

thts wrong

OpenStudy (anonymous):

Oop, sorry! sum = 0 + 1 + 2 = 3 => x = 450/ 3 = 150

OpenStudy (anonymous):

wht abt this one: \[\sum_{x=3}^{5} (x-3)x =90\]

OpenStudy (anonymous):

x = 90 / 3 = 30

OpenStudy (anonymous):

are u positive?

OpenStudy (anonymous):

yes, not hurry this time!

OpenStudy (anonymous):

100% sure?????

OpenStudy (anonymous):

Sum = ( 3 - 3 ) + ( 4 - 3 ) + ( 5 -3) = 0 + 1 + 2 = 3

OpenStudy (anonymous):

can u help me with another question?

OpenStudy (anonymous):

Okie!

OpenStudy (anonymous):

nevermind i got it. thanks for ur help

OpenStudy (anonymous):

:)